Q. The sum of the magnitudes of two forces acting at a point is 18 and the magnitude of their resultant is 12. if the resultant is at 90degree with the force of smaller magnitude , what are the , magnitude of the forces?


a)12,5            b)14,4              c)5,13                  d)10,8

Grade:11

1 Answers

abhishek athreya
13 years ago

let the forces be F1 and F2 

F1+F2=18----------(1)

let the angle between them be 'a'

let their resultant be R

F1^2+F2^2+2FI*F2*cos(a)=R^2--------------------(2)

also 

tan90=F2sin(a)/F1+F2cos(a)------------------------(3)

 

solve (1) (2) and (3)

u will get F1 = 5

and F2=13

 

hope it helped